OpenStudy (nuccioreggie):

A closed container has 6.08 ⋅ 1023 atoms of a gas. Each atom of the gas weighs 1.67 ⋅ 10–24 grams. Which of the following shows and explains the approximate total mass, in grams, of all the atoms of the gas in the container? 1.02 grams, because (6.08 ⋅ 1.67) ⋅ (1023 ⋅ 10-24 ) = 10.1536 ⋅ 10–1 0.78 grams, because (6.08 + 1.67) ⋅(1023 ⋅ 10–24 ) = 7.75 ⋅ 10–1 7.75 grams, because (6.08 + 1.67) ⋅ (1023 ⋅ 10–24 ) = 7.75 10.15 grams, because (6.08 ⋅ 1.67) ⋅ (1023 ⋅ 10–24 ) = 10.1536

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (michele_laino):

total weight W is: \[W = 6.08 \times {10^{23}} \times 1.67 \times {10^{24}}\]

OpenStudy (michele_laino):

oops... \[\Large W = 6.08 \times {10^{23}} \times 1.67 \times {10^{ - 24}}\]

OpenStudy (nuccioreggie):

What thso its B

OpenStudy (michele_laino):

are you sure?

OpenStudy (nuccioreggie):

Im asking

OpenStudy (nuccioreggie):

Do you know the answer choice

OpenStudy (michele_laino):

hint: 6.08*1.67= 10.1536

OpenStudy (nuccioreggie):

D

OpenStudy (michele_laino):

hint: 10^23* 10^(-24)=10^(-1)

OpenStudy (nuccioreggie):

A

OpenStudy (michele_laino):

that's right!